### v1

# Kolmogorov-Arnold Networks: Approximation and Learning Guarantees for Functions and their Derivatives

Abstract
Inspired by the Kolmogorov–Arnold superposition theorem, Kolmogorov–Arnold Networks (KANs) have recently emerged as an improved backbone for most deep learning frameworks, promising more adaptivity than their multilayer perception (MLP) predecessor by allowing for trainable spline-based activation functions. In this paper, we probe the theoretical foundations of the KAN architecture by showing that it can optimally approximate any Besov function in $B^{s}_{p,q}(\mathcal{X})$ on a bounded open, or even fractal, domain $\mathcal{X}$ in $\mathbb{R}^{d}$ at the optimal approximation rate with respect to any weaker Besov norm $B^{\alpha}_{p,q}(\mathcal{X})$; where $\alpha<s$. We complement our approximation guarantee with a dimension-free estimate on the sample complexity of a residual KAN model when learning a function of Besov regularity from $N$ i.i.d. noiseless samples. Our KAN architecture incorporates contemporary deep learning wisdom by leveraging residual/skip connections between layers.
1 Introduction
Many contemporary deep learning backbones leverage trainable activation functions, evolving from their traditional multi-layer perception (MLP) predecessors, to achieve superior adaptivity or training stability. Standard examples range from the SwiGLU [40] activation in transformer networks to the adaptive activations in Kolmogorov-Arnold networks [29] typically realized as a linear combination of B-splines, see e.g. [12], and some interesting theoretical thought experiments are nested neural networks [53]. In this paper, we focus on models of the latter type with an additional residual connection between layers, introduced in the ResNet architecture [18], guided by contemporary deep learning wisdom which recognizes the practical benefits of skip connection and has seen a wide-spread incorporation in contemporary deep learning from GNNs [8] and transformers [44]. Residual connections are theoretically founded and known to positively regularize a deep learning model’s loss landscape [36], they allow for narrower networks to maintain their universality [28] as compared to networks without skip connections [35, 19], and they often have no negative drawback on their approximation rates of those models [17]. We cal these KANs augmented by a residual connection Res-KANs.
Much of the theoretical support for KANs is rooted in the Kolmogorov-Arnold representation theorem; see e.g. [24] for an optimized formulation. These results either provide approximation guarantees for functions of a composition form [30], similarly to the composition sparsity [34, 10] literature, they offer approximation guarantees in the $L^{p}$-norms [46] by encoding ReLUk-neural network structure and subsequently transferring their $L^{p}$-type approximation guarantees [7, 48, 16, 32] or they show that there exist activation functions which allow them to realize a Kolmogorov-Arnold-type superposition representation of any function [34] which is reminiscent of the super-expressive activation function literature [50, 52, 22, 45].
Though these results all support the expressive power of the KAN paradigm, their applicability to the partial differential equation-type (PDE) problems, e.g. arising in physics, engineering, biology, optimal control, or finance, can still be limited. This is because, for PDE applications, one needs to approximate the function itself but typically converge the AI model toward target functions’ higher-order derivatives as well. Additionally, since most KANs are built on splines and spline-type wavelets are known to characterize specific Besov spaces, see e.g. [14, 11] then it seems extremely natural to study these deep learning models in their natural Besov spaces.
Contribution: This paper precisely proves that Res-KANs are efficient approximators of functions on Besov spaces, over any reasonable bounded Lipschitz, or even on any compact fractal domain in $\mathbb{R}^{d}$ (Theorem 1). Importantly, with downstream PDE applications in mind, our approximation guarantees are not (only) in the uniform or $L^{p}$ type but in higher-order Besov norms, which are just arbitrarily weaker than the target function’s regularity.
We complement our main approximation result with an agnostic PAC-learnability guarantee (Theorem 2) showing that KANs can, indeed, learning and Besov function of high-regularity from noiseless training data. Moreover, the sample complexity is not cursed by dimensionality in the high-smoothness regime.
Organization
Our paper is organized as follows. Section 2 overviews the necessary background required to formulate our main results; this includes background on cardinal B-splines, Res-KANs, and on Besov spaces over bounded open Lipschitz domains, and on Ahlfors-regular fractal domains, in $\mathbb{R}^{d}$. Section 3 contains our main results, namely our two approximation theorem (Theorem 1) and our realizable PAC-learning guarantee (Theorem 2). All proofs and additional background, e.g. fat shattering and pseudo-dimension, are only needed for proof details and are relegated to our appendices.
2 Background
We now present the necessary background to formulate our main result.
2.1 Cardinal B-Splines
Kolmogorov-Arnold Networks (KANs) extend the multilayer perception (MLP) architecture by allowing its otherwise fixed and, thus, rigid, univariate activation functions to the individual local structures of the function being approximated. This adaptivity is realized by replacing each activation function by a cardinal B-spline $\mathcal{N}_{k}$ of order $k$ for appropriately learned $k$.
First Cardinal $B$-Splines $I$ $\mathcal{N}_{I}$ 00 $\mathbf{1}_{[0,1)}(x)$ $1$ $\operatorname{ReLU}(x)-2\operatorname{ReLU}(x{-}1)+\operatorname{ReLU}(x{-}2)$ $2$ $\frac{\operatorname{ReLU}(x)^{2}}{2}-\frac{3\operatorname{ReLU}(x{-}1)^{2}}{2}+\frac{3\operatorname{ReLU}(x{-}2)^{2}}{2}-\frac{\operatorname{ReLU}(x{-}3)^{2}}{2}$
Figure 1: The cardinal $B$-splines of orders $I=0,1$, and $2$.
As shown, for example, in [33, Equation (4.28)], for any $I\in\mathbb{N}_{+}$, the cardinal B-spline of order $I$ with knots on $10,\dots,I+$ 1 is given by
$\mathcal{N}_{I}(x)=\sum_{j=0}^{I+1}\frac{(-1)^{j}\binom{I+1}{j}}{I!}\,\operatorname{ReLU}(x-j)^{I}$ (1)
for each $x\in\mathbb{R}$. Figure 1 illustrates the cardinal B-splines with $I=1,2$, and $3$; detailed in the following examples.
2.2 Residual Kolmogorov-Arnold Networks (Res-KANs)
The core idea behind the KAN is to allow the activation function to be trainable, where we allow mixtures of different spline degrees via sending any $x\in\mathbb{R}$ to
$\sigma_{\beta:I}(x)=\sum_{i=0}^{I}\beta_{i}\,\mathcal{N}_{i}(x)$ (2)
where the trainable parameter $\beta\in\mathbb{R}^{I+1}$. In a KAN, each activation function acts componentwise but with trainable parameter depending on the neuron it is activating, that is: for each $d\in\mathbb{N}_{+}$, every $x\in\mathbb{R}^{d}$, and every $\mathbf{\beta}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\beta_{1},\dots,\beta_{k})\in\mathbb{R}^{(I+1)\times d}$ we have
$\sigma_{\mathbf{\beta}:I}\bullet(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{(}\sigma_{\beta_{j}:I}(x_{j})\big{)}_{j=1}^{d}.$ (3)
We now introduce the core of our residual KAN networks, which ensure that no signal is lost during activation by incorporating an additional residual connection, which also have become standard in modern AI as they additionally stabilize training dynamics by preserving gradient flow by regularizing the neural network’s loss landscape [36], and avoid vanishing gradient problems which can be caused by normalizing layers. As in [1], we allow first any residual connections to be skipped or focused on, if need be via a trainable gating mechanism
$\mathcal{L}(x|A,b,\beta,G:I)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\underbrace{\sigma_{\mathbf{\beta}:I}\bullet(Ax+b)}_{\text{KAN Layer}}+\underbrace{Gx}_{\text{Residual Connection}}$ (4)
here the layer input and output dimensions respectively as $d_{in},d_{out}\in\mathbb{N}_{+}$ $A,G$ are $d_{out}\times d_{in}$-matrices with a matrix $G$ diagonal; meaning $G_{i,j}=0$ if $i\neq j$ (not $G$ need not be a square matrix), $\beta$ is a $(I+1)\times d_{\ell+1}$ matrix, $b\in\mathbb{R}^{d_{\ell+1}}$.
Though the composition of these KAN layers does define a meaningful function, that object may not be very regular, in the sense that it may have few higher-order derivatives; making it problematic for PDE applications wherein a high degree of regularity is required. There are two solutions: 1) the $\beta_{i}$ coefficients are all zero small $i$, or 2) we incorporate a simple smoothing layer at the output. We opt for the section option, as then any function implemented by our smoothed residual KANs is necessary smooth; i.e. infinitely differentiable.
Definition 1 (Residual KANs (Res-KANs)).
Let $d,D,I\in\mathbb{N}_{+}$ and $\alpha>0$. A residual Kolmogorov-Arnold network (Res-KAN) is a function $\hat{f}:\mathbb{R}^{d}\to\mathbb{R}^{D}$ with representation
$\hat{f}$ $=A^{(L)}f^{(L)}+b^{(L)}$ (5)
$f^{(l)}$ $=\mathcal{L}(f^{(l-1)}|A^{(l)},b^{(l)},\beta^{(l)},G^{(l)}:I)\mbox{ for }l=1,\dots,L$
$f^{(0)}(x)$ $=x$
where, for $l=1,\dots,L$, $A^{(l)}$ and $G^{(l)}$ are $d_{l+1}\times d_{l}$ matrices with $G$ diagonal, $\beta$ is a $(1+I)\times d_{l+1}$ matrix, $b\in\mathbb{R}^{d_{l+1}}$, $d_{0},\dots,d_{L+1}\in\mathbb{N}_{+}$ with $d_{0}=d$ and $d_{L+1}=D$; furthermore, $\beta^{(l)}$ satisfies the sparsity pattern (ensuring smoothness)
$\beta^{(l)}_{i,j}=0\mbox{ for }i<\lceil\alpha\rceil.$ (6)
We denote the class of all Res-KANs with $L$ hidden layers, width $W\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\max_{l=1,\dots,L+1}\,d^{(l)}$, adaptivity parameter $I$, and smoothness parameter $\alpha$ by $\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R}^{D})$.
2.3 Besov Spaces
We begin with the definition of the Besov spaces on any non-empty open domain $\Omega$ in $\mathbb{R}^{d}$.
Besov Spaces on Euclidean Domains
Let $0<p\leq\infty$, and $0<\alpha$. The modulus of smoothness of order $\alpha$ of an $f\in L^{p}(\Omega)$, defined with respect to the restricted Lebesgue measure on $\Omega$ is defined for each $t>0$
$\omega_{\alpha}(f,t)_{p}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sup_{\delta\in\mathbb{R}^{d}\,0<\|\delta\|\leq t}\,\|\Delta_{\delta}^{\lceil\alpha\rceil}(f,\cdot)\|_{L^{p}(\lceil\alpha\rceil h)}$
[…]
We will focus on the following class of domains: An open set $\Omega$ is called an extension domain in $\mathbb{R}^{d}$ if: there is some $\epsilon,\delta>0$ such that for any $x,y\in\Omega$ satisfying
$\|x-y\|\leq\delta$
there exists a rectifiable curve $\gamma:[0,1]\to\mathbb{R}^{d}$ of length at-most $C_{0}\|x-y\|$ with $\gamma(0)=x$ and $\gamma(1)=y$ such that for any time $t\in[0,1]$
$\inf_{u\in\partial\Omega}\,\|\gamma(t)-u\|\geq\epsilon\min\{\|\gamma(t)-x\|,\|\gamma(t)-y\|\}.$
We are interested in extension domains due to the extendability of functions therein to all of $\mathbb{R}^{d}$, while also being able to ensure that their Sobolev regularity is preserved under these extensions; see e.g. [23, 9]. Extension domains, in the above sense, are often referred to as $(\epsilon,\delta)$-domains in harmonic analysis. However, we avoid this terminology here to prevent confusion with its standard use in approximation and learning theory, where $\epsilon$ typically denotes an approximation error and $\delta$ the probability that the approximation is valid.
Besov Spaces on Fractals
We begin with the definition of a Besov space on an arbitrary well-behaved, possibly fractional, domains $\mathcal{X}$ in some Euclidean space $\mathbb{R}^{d}$ as introduced in [20]. Let $0<n\leq d$ and let $\mathcal{H}^{n}$ denote the $n$-dimensional Hausdorff (outer) measure on $\mathbb{R}^{d}$. We denote a closed $\ell^{\infty}$-ball of radius $r>0$ centered at some $x\in\mathbb{R}^{n}$ by $Q(x,r)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{y\in\mathbb{R}^{d}:\,\|x-y\|_{\infty}\leq r\}=[x-r/2,x+r/2]^{d}$. A subset $\mathcal{X}\subseteq\mathbb{R}^{d}$ is called Ahlfors $n$-regular if: there are constants $0<c\leq C$ such that for all $0<r\leq\operatorname{diam}(\mathcal{X})$ and each $x\in\mathcal{X}$
$cr^{n}\leq\mathcal{H}^{n}\big{(}\mathcal{X}\cap Q(x,r)\big{)}\leq Cr^{n}.$ (8)
For $1\leq p<\infty$, we define the normalized local best polynomial approximation energy as
[…]
where $\mathcal{Q}=\operatorname{Ball}_{\ell^{\infty}}(x,r)$ for some $x\in\mathcal{X}$ and some $r>0$, $\mathbb{R}_{k-1}[x_{1},\dots,x_{d}]$ is the vector space of polynomials on $x_{1},\dots,x_{d}$ of degree at-most $k-1$ with the convention that $\mathbb{R}_{-1}[x_{1},\dots,x_{d}]\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{0\}$ is the trivial vector space.
Definition 2 (Besov Space on an Ahlfors-Regular Sets).
Let $0<n\leq d$, $\mathcal{X}\subseteq\mathbb{R}^{d}$ be Ahlfors $n$-regular, and let $\alpha>0$, $1\leq p,q\leq\infty$. The Besov space $B^{\alpha}_{p,q}(\mathcal{X})$ consists of all functions $f\in L^{p}(\mathcal{X})$ for which the norm
$\|f\|_{B^{\alpha}_{p,q}(\mathcal{X})}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\|f\|_{L^{p}(\mathcal{X})}+\left(\int_{0}^{1}\left(\frac{\|\mathcal{E}_{k}(f,Q(\cdot,\tau))\|_{L^{u}(\mathcal{X})}}{\tau^{\alpha}}\right)^{q}\frac{d\tau}{\tau}\right)^{\frac{1}{q}},$
is finite; where $1\leq u\leq p$ and $k$ is an integer such that $\alpha<k$.
Importantly, by [21, Theorem 3.6], the definition of the Besov space $B^{\alpha}_{p,q}(\mathcal{X})$ above does not depend on the (arbitrary) choice of parameters $k$ and $u$; granted that $k>\alpha$ and $1\leq u\leq p$.
3 Main Results
We are now in place to state our main approximation guarantee.
Theorem 1 (Approximation Guarantees for Res-KANs in Besov Norm).
Let $0<\alpha<s<\infty$ and $0<p,q<\infty$, $d-1<n<d$ and $\mathcal{X}\subseteq[0,1]^{d}$ either be: (1) an $(\epsilon,\delta)$-domain or (2) Ahlfors $n$-regular for some $d-1<n<d$ and additionally $1\leq p,q$. In case (1) set $\alpha^{\star}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\alpha$ and in case (2) set $\alpha^{\star}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\alpha-(n-d)/p$. For every $f\in B_{p,q}^{\alpha^{\star}}(\mathcal{X})$ and every “simultaneous approximation error” $\varepsilon>0$ there exists a Res-KAN $\hat{f}:\mathbb{R}^{d}\to\mathbb{R}$ such that $\hat{f}\in B_{p,q}^{r}(\mathcal{X})$ satisfying
$\|f-\hat{f}|_{\mathcal{X}}\|_{B_{p,q}^{r}(\mathcal{X}))}<\varepsilon.$
Moreover, $\hat{f}$ has width $\mathcal{O}(\varepsilon^{1/((\alpha^{\star}-s)q)})$, depth at-most $d+1$, and at-most $\mathcal{O}\big{(}d^{2}\varepsilon^{1/((\alpha^{\star}-s)q)}\big{)}$ from Lemma 4 non-zero parameters.
Proof.
See Appendix A. ∎
We now complement our approximation guarantee for residual KANs with our learning guarantee. We operate within the noiseless PAC-learning framework, which generalizes the realizable PAC-learning setup (see e.g. [39, 4]) as it most closely aligns with our approximation guarantee. That is, we ask, given any target function of a given Besov-regularity $f\in B_{p,q}^{\alpha}(\mathcal{X})$, how many i.i.d. samples $N$ are required to ensure that the in-sample performance of any residual Kolmogorov-Arnold network $\hat{f}$ closely mirrors its out-of-sample performance?
The realizable setting, most closely mirrors the approximation setting, as it assumes that we are directly observing paired samples of input and output data, uncorrupted by exogenous measurement noise. This is akin to the approximation theorems consider unlimited amounts of uncorrupted paired samples. We quantify the true error between any given Res-KAN $\hat{f}$ and any target $f$ by their true risk $\mathcal{R}_{\mathbb{P}}(f|\hat{f})$ and their empirical risk ^$\hat{\mathcal{R}}_{\mathbb{P}}^{N}(f|\hat{f})$ defined by
$\mathcal{R}_{\mathbb{P}}(f|\hat{f})=\mathbb{E}_{X\sim\mathbb{P}_{X}}\big{[}\|\hat{f}(X)-f(X)\|\big{]}\mbox{ and }\hat{\mathcal{R}}_{\mathbb{P}}^{N}(f|\hat{f})=\frac{1}{N}\sum_{n=1}^{N}\|\hat{f}(X_{n})-f(X_{n})\|.$
Under the simplifying assumption that the target function is of unit Besov norm, we have.
Theorem 2 (PAC-Learning Guarantees for Res-KANs: With Noiseless Besov Data).
Suppose that $\mathcal{X}$ is a Lipschitz domain, let $1\leq\tau\leq\infty$, $1\leq p,q<\infty$ and $\alpha>\left(d\left(1/p-1/\tau\right)\right)_{+}$ and let $L,I,W\in\mathbb{N}_{+}$. For every probability measure $\mathbb{P}_{X}\in\mathcal{P}(\mathcal{X})$ each training set of i.i.d. samples $X_{1},\dots,X_{N}\sim\mathbb{P}_{X}$ every approximation error $\varepsilon>0$ and every failure probability $0<\delta\leq 1$ if
$N\in\mathcal{O}\Bigl{(}\epsilon^{-2-d/\alpha}\,(\ln(1/\varepsilon))^{2}+\epsilon^{-2}\ln(1/\delta)\Bigr{)}.$
then
$\mathbb{P}\Biggl{(}\sup_{f,\hat{f}}\,\Big{|}\mathcal{R}_{\mathbb{P}}(f|\hat{f})-\hat{\mathcal{R}}_{\mathbb{P}}^{N}(f|\hat{f})\Big{|}\leq\varepsilon\Biggr{)}\geq 1-e^{-cN\epsilon^{2}+d^{\star}\ln^{2}\Big{(}\tfrac{d^{\star}}{\epsilon}\Big{)}}$
where the supremum is taken over all $\hat{f}\in\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})$ and all $f\in B_{p,q}^{\alpha}(\mathcal{X})$ with Besov norm $\|f\|_{B_{p,q}^{\alpha}(\mathcal{X})}\leq 1$.
Proof.
See Appendix B. ∎
3.1 Sanity Checks
Since the original KAN paper [30] and its many variants [47, 54, 37] are now well-established as effective learners, we only perform brief sanity checks to validate our theory and confirm that the addition of the residual connection has a consistently positive impact on the KAN architecture. We first then verify that one can indeed learn a function in the Besov space $B^{\alpha}_{2,2}([0,1])$, but not in $B^{\alpha+1}_{2,2}([0,1])$, as well as its derivatives. We then verified that the traditional KAN build and our mild variant with residual connections both offer similar performance. We verify that the training loss converges relatively steadily during training.
A Non-Smooth Besov Function of a Specific Regularity
The Besov space $B^{\alpha}_{2,2}([0,1])$ coincides with the Sobolev space $H^{\alpha}([0,1])$ for integer $\alpha>0$. Thus, we may briefly validate our theoretical results for the family of functions $\{f_{\alpha}\}_{\alpha=1}^{10}$, where for each such $\alpha$ we set
$f_{\alpha}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\frac{x^{\alpha}\log(x)}{\alpha!}.$ (10)
We chose this example since, $f_{\alpha}$ is a classical example of a function belonging to $H^{\alpha}([0,1])$ but not to $H^{\alpha+1}([0,1])$. To see this, notice that $f_{\alpha}$ is $\alpha$-times differentiable, not only weakly, with $s^{th}$ derivative bounded below near 00 by
$\frac{d^{{\alpha}+1}}{dx^{{\alpha}+1}}x^{\alpha}\log(x)\succsim\frac{1}{x}$
therefore $\lim\limits_{x\downarrow 0}\frac{d^{{\alpha}+1}}{dx^{{\alpha}+1}}x^{\alpha}\log(x)=\infty$. We have chosen the normalization factor of $1/{\alpha!}$ since it is the leading coefficient of the ${\alpha-1}^{rst}$ derivative of $x^{\alpha}\log(x)$. Thus, the Sobolev/Besov $\|\cdot\|_{B^{\alpha}_{2,2}([0,1])}$ norm of each $f_{\alpha}$ remains roughly on the same scale (about unity) and are indeed comparing apples-to-apples between derivative levels of $\alpha$.
We Do Actually Learn A Function And Its Derivatives?
Next, we verify Theorems 1 and 2 are indeed reflected in practice; namely, that we can both approximate function and its higher (weak) derivatives from training data. By the Pointcaré-inequality, it suffices to compare the mean squared error (MSE) between $s^{th}$ derivative our prediction and the target data on $[0,1]$. Iterating the Pointcaré-inequality implies that we have control over the function itself and all the first $s-1$ derivatives upon controlling our MSE to the $s^{th}$ derivative; i.e.
$\|f\|_{L^{2}([0,1])}\lesssim\dots\lesssim\Big{\|}\frac{\partial^{s}f}{\partial x^{s}}\Big{\|}_{L^{2}([0,1])}\lesssim\Big{\|}\frac{\partial^{s+1}f}{\partial x^{s+1}}\Big{\|}_{L^{2}([0,1])}.$
Note, in practice, all derivatives are computed numerically via autograd. Thus, we train on the MSE loss augmented with the MSE between the ${\alpha}^{th}$ derivative of our Res-KAN and the target function, i.e. for a hyperparameter $0\leq\lambda\leq 1$ controlling our focus on higher derivatives during training, we optimize the following loss function
$\operatorname{loss}_{\lambda}(\hat{f})\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\frac{(1-\lambda)}{N}\sum_{n=1}^{N}\big{(}f(X_{n})-\hat{f}(X_{n})\big{)}^{2}+\frac{\lambda}{N}\sum_{n=1}^{N}\Big{(}\frac{\partial^{s}f}{\partial x^{s}}(X_{n})-\frac{\partial^{s}\hat{f}}{\partial x^{s}}(X_{n})\Big{)}^{2}.$
We allow $\lambda$ to linearly decrease from 1 to 0 during training, so that the KAN initially learns the overall function structure (encoded in its derivatives) before fine-tuning its pointwise value approximation in the final SGD iterations.
Figure 2: Evolution of training loss, test prediction, and derivative approximation quality when learning the functions $f_{\alpha}$ in (10) for $\alpha=3$ and $\alpha=5$.
Do The Residual Connection Adversely Impact the KAN Build?
We now compare the KAN build with, and without, the additional residual connection. Our objective is simply to perform a quick check validating standard deep learning wisdom, that its incorporation, can only benefit the model. Each model has width $200$, two hidden layers, and is trained in-sample with $10^{3}$ datapoints uniformly sampled on $[-2,2]$ then tested, mimicking the realizable setting of Theorem 2, on a grid of $10^{2}$ test points in $[-2,2]$. The experiment confirms that the KAN builds can capture the low-regularity structures, e.g. the spike and rapid a-periodic osculations, and both builds offer similar performance. The results of our experiments are plotted as a function of the integer values of $\alpha$ from $1$ to $S=30$ in Figure 3. As we see, both KAN builds with or without residual connection offer similar predictive performance regardless of the degree of Sobolev-Besov regularity available in the target function.
Figure 3: Both the KAN and Res-KAN build offer similar approximation efficacy across different Sobolev-Besov smoothness levels.
Functions Beyond our Guarantees
Lastly, we briefly verify that the ability of the KAN with residual can handle pathological regression tasks. Figure 4 briefly considers two such cases when the target function exhibits a rapid a-periodic osculatory, e.g. $1/\cos(x)$, or when it exhibits sharp cusps, e.g. $1/\sqrt[10]{x}$. As we see, the training loss descents, the challenging pattern can be learned to reasonable accuracy, and the higher derivatives of the Res-KAN indeed do also converge to those of the pathological target function.
Figure 4: Two challenging one-dimensional functions to learn - using $\operatorname{ReLU}$ and $\operatorname{ReLU}^{2}$ MLPs, as well as the standard and the residual KAN builds.
A detailed numerical investigation confirms that, for the relevant class of Besov-type functions and beyond, including a residual connection in the KAN architecture does not negatively impact performance. Consequently, practitioners who prefer implementing the standard, non-residual KAN can expect our results to transfer seamlessly to this simpler setting. Nevertheless, the residual connection may offer practical benefits, particularly in more complex regimes, by enhancing optimization stability and facilitating convergence in deeper or wider network configurations.
4 Conclusion
In this paper, we establish the theoretical foundations of Kolmogorov–Arnold Networks (KANs), showing that they can approximate any Besov function on a bounded or even fractal domain at the optimal rate in Theorem 1. We also provide a dimension-free sample complexity bound for learning such functions with a residual KAN model in Theorem 2. Due to the deep connection between Besov spaces and splines [14, 11] we believe that this is a very natural setting to quantify the power of Kolmogorov-Arnold networks. Our results relied on a KAN build which incorporated residual connections, aligning with modern deep learning practices. Simple toy experiments further confirm that adding residual connections does not degrade performance, and the KAN retains accuracy similar to that of its non-residual counterpart.
Acknowledgements
A. Kratsios acknowledges financial support from an NSERC Discovery Grant No. RGPIN-2023-04482 and No. DGECR-2023-00230. They both also acknowledge that resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute [^1] https://vectorinstitute.ai/partnerships/current-partners/.
T. Furuya was supported by JSPS KAKENHI Grant Number JP24K16949, 25H01453, JST CREST JPMJCR24Q5, JST ASPIRE JPMJAP2329.
Appendix A Proof of Theorem 1
We now prove Theorem 1. We first verify that residual KANs can locally-implement the multiplication operation. The next lemma serves as an exact version of the well-known fact that: $\tanh$-MLPs [13], ReLU MLPs [49], and more generally MLPs with smooth activation function [26, 27, 51], can approximately implement the $d$-fold multiplication operation on arbitrarily large hypercubes. This next result shows that residual KANs can exactly implement the $d$-fold multiplication operation, locally, on arbitrarily large hypercubes.
Lemma 1 (Exact Multiplication on Arbitrarily-Large Hypercubes).
For every $d\in\mathbb{N}_{+}$ and each $M>0$ there exist a Res-KAN ×$\times^{2}_{d}:\mathbb{R}^{d}\to\mathbb{R}$ satisfying: for each $x\in[-M,M]^{d}$
×$\times^{2}_{d}(x)=\prod_{i=1}^{d}\,x_{i}$
Moreover ×$\times^{2}_{d}$ has depth $d$, width at-most $12d+$ 1, and at-most $\frac{5d^{2}+21d}{2}$ non-zero parameters.
Proof.
Step 00 - Implementing The $n$-fold Square:
If $\lceil\alpha\rceil\geq 2$ then $I+1\geq 3$ and may legitimately set $\beta\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\mathbf{0}_{d},\mathbf{0}_{d},\mathbf{1}_{d},\mathbf{0}_{d},\dots,\mathbf{0}_{d})$ be the $(I+1)\times d$ matrix with all entries zero except with its third column populated only by $1$s. Consider the layer
[…]
Since the knots of each $N_{i}$ are on the integer points $10,\dots,I+$ 1 then the restriction of the rescaled B-splines $\mathcal{N}_{2}(2Mx-M)$ to $[-M,M]$ is simply
$\hat{m}_{M}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathcal{N}_{2}(2Mx-M)|_{[-M,M]}=\frac{\binom{I+1}{j}}{I!}\operatorname{ReLU}(x)^{2}|_{[-M,M]}.$ (12)
By construction, $\hat{m}_{M}$ is defined by at most $14n+$ 1 non-zero parameters.
Step $1$ Implementing The $n$-Multiplication:
As in the proof of [16, Proposition 1], for each $n\in\mathbb{N}_{+}$, consider the partial multiplication map
$\phi_{n+1:\uparrow,\times}:\mathbb{R}^{n+1}$ $\rightarrow\mathbb{R}^{n}$
$(x_{i})_{i=1}^{n+1}$ $\mapsto(x_{1}x_{2})\oplus(x_{i})_{i=3}^{n+1}.$
Again, appealing to the proof of [16, Proposition 1], there matrices
$W_{2}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{(}(2,2,2)\oplus\mathbf{0}_{n-1}|\mathbf{0}_{n-1,n+2}\big{)},\,W_{1}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\begin{pmatrix}1/2&1/2\\ -1/2&0\\ 0&-1/2\end{pmatrix}\oplus\mathbf{0}_{n-1},\mbox{ and }W_{s}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathbf{0}_{1}\oplus I_{n}$
are such that $\phi_{n+1:\uparrow,\times}$ can be globally implemented as
$\phi_{n+1:\uparrow,\times}(x)=W_{2}\operatorname{ReLU}^{2}\big{(}W_{1}x+\mathbf{0}_{n+1}\big{)}+W_{s}x.$ (13)
Now, combining (12) and (13), we have that: for every $M>0$ and each $x\in[-M,M]^{n}$
$\phi_{n+1:\uparrow,\times}(x)=\hat{\phi}_{n+1:\uparrow,\times:M}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}W_{2}\hat{m}_{M}\big{(}W_{1}x+\mathbf{0}_{n+1}\big{)}+W_{s}x.$ (14)
Moreover, $\hat{\phi}_{n+1:\uparrow,\times:M}$ is defined by at-most $8+5n$ non-zero parameters. Now, iteratively appealing to our above construction for $n=2,\dots,d$, we find that: for each $x\in[-M,M]^{d}$
$\prod_{i=1}^{d}\,x_{i}=\phi_{2:\uparrow,\times}\circ\cdots\circ\phi_{d:\uparrow,\times}(x)=\hat{\phi}_{2:\uparrow,\times:M}\circ\cdots\circ\hat{\phi}_{d:\uparrow,\times:M}(x)$ (15)
[…]
Together, we may approximately implement a basic component of a spline-type multi-resolution analysis (MRA), see [31], type of characterization of the Besov space $B^{\alpha}_{p,q}([0,1]^{d})$ derived in [15]. Returning to our cardinal B-splines above, expressed in (1), we consider the $d$-fold tensor product of such splines $\mathcal{N}_{d:I}:\mathbb{R}\to\mathbb{R}$ given for each $x\in\mathbb{R}^{d}$ by
$\mathcal{N}_{d:I}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\prod_{k=1}^{d}\,\mathcal{N}_{I}(x_{k}).$ (16)
Next, for each $k\in\mathbb{N}_{+}$ let $\mathbb{D}_{k}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{\prod_{i=1}^{d}\,[x_{i}-2^{k-1},x_{i}+2^{k-1}]:\,2^{k}x\in\mathbb{Z}^{d}\}$ denote the set of dyadic cubes in $\mathbb{R}^{d}$ with side-length $2^{-k}$ centred at points in the dyadic lattice $2^{k}\mathbb{Z}^{d}$. For any $\Omega\subseteq\mathbb{R}^{d}$, we write $\mathbb{D}_{k}^{\Omega}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{Q\in\mathbb{D}_{k}:\,Q\cap\Omega\neq\emptyset\}$. We consider a spline-based MRA based on (16) over the cube $[0,1]^{d}$ is constructed as follows: for every $k\in\mathbb{N}_{+}$ and each dyadic cube $Q\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}Q_{\mathbf{j:k}}\in\mathbb{D}_{k}$ we associate a spline $\mathcal{N}_{\mathbf{j},k:I}:\mathbb{R}\to\mathbb{R}$ defined by
$\mathcal{N}_{\mathbf{j},k:I}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathcal{N}_{I}(2^{k}x-\mathbf{j}).$ (17)
Our next lemma shows that the Res-KAN networks can exactly emulate the basic wavelet-type splines in (17).
Lemma 2 (Multi-Dimensional Cardinal $B$-Spline Implementation).
Let $I,d,k\in\mathbb{N}_{+}$, and $\mathbf{j}\in 2^{-k}\mathbb{Z}^{k}$. There exists a Res-KAN network $\hat{\mathcal{N}}_{d:I}:\mathbb{R}^{d}\to\mathbb{R}$ such that: for each $x\in\mathbb{R}^{d}$
$\hat{\mathcal{N}}_{k,\mathbf{j}:I}(x)=\mathcal{N}_{k,\mathbf{j}:I}(x)$
Moreover, $\hat{\mathcal{N}}_{k,\mathbf{j}:I}$ has width at-most $12d+$ 1, depth $d+1$, and at-most $\frac{5d^{2}+25d}{2}$ non-zero parameters. Furthermore, only the $2d$ non-zero parameters in layer depends on $k$ and on $\mathbf{j}$.
Proof of Lemma 2.
We consider the case where $d>1$ with the case where $d=1$ following the definition of the res-KAN. Consider the Res-KAN $\hat{\mathcal{N}}_{d:I}:\mathbb{R}^{d}\to\mathbb{R}$ given by for each $x\in\mathbb{R}^{d}$
$\hat{\mathcal{N}}_{d:I}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\times^{2}_{d}\Big{(}\sigma_{e_{I+1}}\bullet I_{d}x+\mathbf{0}_{d}\Big{)}$
where $e_{1},\dots,e_{I+1}$ are the standard basis vectors of $\mathbb{R}^{I+1}$, $\bullet$ denotes componentwise composition, and ×$\times_{d}^{2}$ is as in Lemma 1. By Lemma 1 $\hat{\mathcal{N}}_{d:I}$ has width at-most $12d+$ 1, depth $d+1$, and at-most $\frac{5d^{2}+23d}{2}$ non-zero parameters; moreover, we have
$\hat{\mathcal{N}}_{d:I}(x)$ $=\prod_{i=1}^{d}\Big{(}\sigma_{e_{I+1}}\bullet I_{d}x+\mathbf{0}_{d}\Big{)}$
$=\prod_{i=1}^{d}\Big{(}\mathcal{N}_{I}(x_{k})\Big{)}$
$=\mathcal{N}_{d:I}(x)$
where the last equality holds by definition of $\mathcal{N}_{d:I}$ in (16). Consequently, for each $k\in\mathbb{N}_{+}$ and each $\mathbf{j}\in 2^{-k}\,\mathbb{Z}^{k}$ we have
$\hat{\mathcal{N}}_{\mathbf{j},k:I}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathcal{N}_{d:I}(2^{k}x-\mathbf{j})=\hat{\mathcal{N}}_{d:I}(2^{k}x-\mathbf{j}).$
Since the map $x\mapsto 2^{k}x-\mathbf{j}$ is affine map from $\mathbb{R}^{d}$ to itself, then $\hat{\mathcal{N}}_{k,\mathbf{j}:I}(\cdot)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\hat{\mathcal{N}}_{d:I}(2^{k}\cdot-\mathbf{j})$ also has depth $d$, width at-most $12d+$ 1, and simple count show that it has at-most $\frac{5d^{2}5d}{2}$ non-zero parameters. ∎
Using one of the main results of [15], we are now able to describe the Besov spaces $B_{p,q}^{\alpha}([0,1]^{d})$ in terms of the Res-KAN networks. We reiterate that the key point here is the approximability of the Besov norm.
Lemma 3 (Approximation in Besov Norm by Normalized Residual KAN Network).
Let $0<q,p\leq\infty$ and $\alpha<s<\infty$, for each $f\in B^{s}_{p,q}([0,1]^{d})$. For every “simultaneous approximation error” $\varepsilon>0$, there exists a Res-KAN $\hat{f}_{\varepsilon}:\mathbb{R}^{d}\to\mathbb{R}$ satisfying
$\big{\|}f-\hat{f}_{\varepsilon}\big{\|}_{B^{\alpha}_{p,q}([0,1]^{d})}<\varepsilon.$
Moreover, $\hat{f}$ has width at-most $(2d+1)\big{(}2^{K+1}-2\big{)}$, depth at-most $d+1$, and no more than $\left(\frac{5d^{2}+23d}{2}+1\right)(2^{K+1}-1)$ non-zero parameters; where $K\leq\big{\lceil}\log_{2}\varepsilon^{1/((\alpha-s)q)}-1\big{\rceil}$.
Proof of Lemma 3.
Step $1$ Asymptotic Spline Expansion: Set $r\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\lceil\alpha\rceil$, $\lambda\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\min\{r,r-1+\frac{1}{p}\}$, $0<q,p<\infty$, and $0<\alpha<\lambda$. Set $I\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}r$. By [14, Corollary 5.3] we know that any $f\in L_{p}([0,1]^{d})$ belongs to $B_{p,q}^{s}([0,1]^{d})$ if and only if there is a real sequence $\beta_{\cdot}^{f}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\beta_{Q:k}^{f})_{Q\in\mathbb{D}_{k}^{[0,1]^{d}},k\in\mathbb{N}_{+}}$ such that
$f=\sum_{k\in\mathbb{N}}\sum_{Q_{\mathbf{j}}\in\mathbb{D}_{k}^{[0,1]^{d}}}\beta_{Q:k}^{f}\mathcal{N}_{\mathbf{j},k:r}$ (18)
for all $x\in[0,1]^{d}$ with coefficient sequence $\beta_{\cdot}^{f}$ satisfying
$c\|f\|_{B_{q,p}^{s}([0,1]^{d})}^{q}\leq\|f\|_{\operatorname{spline}:\alpha,p,q}\leq C\|f\|_{B_{q,p}^{s}([0,1]^{d})}^{1}$ (19)
for some absolute constants $0<c\leq C<$ depending only on $p,q,s$, and on $d$; where the quasi-norm $\|\cdot\|_{\operatorname{spline}:\alpha,p,q}$ is given by
$\|f\|_{\operatorname{spline}:\alpha,p,q}^{q}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sum_{k=0}^{\infty}2^{\alpha kq}\,\Biggl{(}\sum_{j\in\Lambda(k)}|\beta_{\mathbf{j},k}|^{p}2^{-kd}\Biggr{)}^{q/p}$ (20)
and where $\Lambda(k)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{\mathbf{j}\in 2^{-k}\mathbb{Z}^{d}\,:\,\beta_{\mathbf{j}}\neq 0\}$ (see [14, page 402 above Equation (4.8)]).
Observe that, again by [14, Corollary 5.3], since $f\in B_{p,q}^{s}([0,1]^{d})$ then $\|f\|_{\operatorname{spline}:s,p,q}<\infty$. Thus, (20) implies that: there is some $C>0$ such that
[…]
Thus, (21) implies the following growth condition on the terms $\Biggl{(}\sum_{j\in\Lambda(k)}|\beta_{\mathbf{j},k}|^{p}2^{-kd}\Biggr{)}^{q/p}$
[…]
Step $2$ Spline Emulation: Applying Lemma 2, once for each $k\in\mathbb{N}_{+}$ and every $\mathbf{j}\in\Lambda(k)$, we deduce the existence of a sequence of Res-KAN networks $\big{\{}\hat{\mathcal{N}}_{k,\mathbf{j}:I}\big{\}}_{k\in\mathbb{N}_{+},\,\mathbf{j}\in\Lambda(j)}$ satisfying the global emulation property
$\hat{\mathcal{N}}_{k,\mathbf{j}:I}(x)=\mathcal{N}_{k,\mathbf{j}:I}(x)$
for each $x\in\mathbb{R}^{d}$. Again, each such $\hat{\mathcal{N}}_{k,\mathbf{j}:I}$ has width at-most $12d+$ 1, depth $d+1$, and at-most $\frac{5d^{2}+25d}{2}$ non-zero parameters and only the first layer depend on $k$ and on $\mathbf{j}$ with all other parameters being shared. Thus, the asymptotic expansion in (18) may be re-expressed purely as a convergent series of Res-KANs
$f=\sum_{k\in\mathbb{N}}\sum_{\mathbf{j}\in\Lambda(k)}\beta_{Q:k}^{f}\hat{\mathcal{N}}_{\mathbf{j},k:I}$ (23)
again, with a (actually the same) coefficient sequence $\beta_{\cdot}^{d}$ inducting a finite quasi-norm (20).
Step $3$ Finite-Parameterized Truncation: We now construct our approximator by truncating the expansion in (23) as follows. For each $K\in\mathbb{N}_{+}$, to be set retroactively, define
$\hat{f}_{K}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sum_{k=0}^{K}\,\sum_{\mathbf{j}\in\Lambda(k)}\beta_{Q:k}^{f}\hat{\mathcal{N}}_{\mathbf{j},k:I}.$ (24)
It remains the bound the Besov norm of $\|f-\hat{f}_{K}\|_{B^{\alpha}_{p,q}([0,1]^{d})}$ above. For this, using (23), (24), and again relying on (2) we have that
$\|f-\hat{f}_{K}\|_{B^{\alpha}_{p,q}([0,1]^{d})}$ $\lesssim\biggl{\|}\sum_{k\in\mathbb{N}}\sum_{\mathbf{j}\in\Lambda(k)}\beta_{Q:k}^{f}\mathcal{N}_{\mathbf{j},k:r}-\sum_{k=0}^{K}\,\sum_{\mathbf{j}\in\Lambda(k)}\beta_{Q:k}^{f}\hat{\mathcal{N}}_{\mathbf{j},k:r}\biggr{\|}_{B^{\alpha}_{p,q}([0,1]^{d})}$
$=\biggl{\|}\sum_{k\in\mathbb{N}}\sum_{\mathbf{j}\in\Lambda(k)}\beta_{Q:k}^{f}\mathcal{N}_{\mathbf{j},k:r}-\sum_{k=0}^{K}\,\sum_{Q_{\mathbf{j}}\in\mathbb{D}_{k}^{[0,1]^{d}}}\beta_{Q:k}^{f}\mathcal{N}_{\mathbf{j},k:r}\biggr{\|}_{B^{\alpha}_{p,q}([0,1]^{d})}$
$=\biggl{\|}\sum_{k=K+1}^{\infty}\sum_{\mathbf{j}\in\Lambda(k)}\beta_{Q:k}^{f}\mathcal{N}_{\mathbf{j},k:r}\biggr{\|}_{B^{\alpha}_{p,q}([0,1]^{d})}$
$\lesssim\sum_{k=K+1}^{\infty}2^{\alpha kq}\,\Biggl{(}\sum_{j\in\Lambda(k)}|\beta_{\mathbf{j},k}|^{p}2^{-kd}\Biggr{)}^{q/p}$
$\lesssim\sum_{k=K+1}^{\infty}C2^{\alpha kq}2^{-skq}$
$=C\sum_{k=K+1}^{\infty}\left(2^{(\alpha-s)q}\right)^{k}$
$=C\frac{2^{(\alpha-s)q(K+1)}}{1-2^{q(\alpha-s)}}$
[…]
where $C_{\alpha,s,q}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}C/(1-2^{q(\alpha-s)}>0$. Thus, for any given $\varepsilon>0$, we retroactively set
$K=\biggl{\lceil}\frac{\log_{2}\varepsilon}{(\alpha-s)q}-1\biggr{\rceil}=\Big{\lceil}\log_{2}\varepsilon^{1/((\alpha-s)q)}-1\Big{\rceil}$ (26)
and re-label $\hat{f}_{\varepsilon}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\hat{f}_{K}$ then, (25) implies that
$\|f-\hat{f}_{\varepsilon}\|_{B^{\alpha}_{p,q}([0,1]^{d})}\lesssim_{\alpha,s,q}\varepsilon$
where $\lesssim_{\alpha,s,q}$ is used to a constant only depending on $\alpha$, $s$, and on $q$.
Step $4$ Verifying Neural Representation: Since, for each $k\in[K]$ and every $j\in\Lambda(k)$, the networks $\hat{\mathcal{N}}_{\mathbf{j},k:I}$ have the same depth then we may represent (24) by a neural network of width
$\sum_{k=0}^{K}\,\sum_{\mathbf{j}\in\Lambda(k)}(2d+1)\leq(2d+1)\sum_{k=1}^{K}2^{k}=(2d+1)\big{(}2^{K+1}-2\big{)}$
depth $d+1$, and with at-most
$\sum_{k=0}^{K}\#\sum_{\mathbf{j}\in\Lambda(k)}\biggl{(}\frac{5d^{2}+23d}{2}+1\biggr{)}\leq\sum_{k=0}^{K}2^{k}\left(\frac{5d^{2}+23d}{2}+1\right)=\left(\frac{5d^{2}+23d}{2}+1\right)(2^{K+1}-1)$
non-zero parameters. Namely via the representation
$\hat{f}_{\varepsilon}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\Biggl{(}\bigoplus_{k\in[K],j\in\Lambda(k)}\beta_{Q:k}^{f}\Biggr{)}^{\top}\Biggl{(}\bigoplus_{k\in[K],j\in\Lambda(k)}\hat{\mathcal{N}}_{\mathbf{j},k:I}\big{(}\mathbf{1}_{\sum_{k\in[K]\sum_{j\in\Lambda(k)}}}\,x\big{)}\Biggr{)}.$
This completes our proof. ∎
We remind the reader that this directly implies Sobolev-norm approximation guarantees since $B_{p,p}^{s}(\Omega)$ is equivalent to the Sobolev space $W^{s,p}(\Omega)$ when $s$ is not an integer; see e.g. [42, Equation and Remark 5.31]. We now extend the conclusion of Lemma 3 to general domains.
A.1 From the Unit Cube to General Domains
We now extend Lemma 3 to general domains. We consider two cases: classical regular domains or domains of fractal type. Both are treated separately, and their joint conclusion is the main result of this section.
A.1.1 Regular Domains
The following result is true on $(\varepsilon,\delta)$-domains in $\mathbb{R}^{d}$. We may now extend our results from the $d$-dimensional unit hypercube to general Lipschitz $(\varepsilon,\delta)$-domains by using Rychkov’s extension operator, introduced in [38], as improved on in [41].
Lemma 4 (Approximation in Besov Norm by Normalized Residual KAN Network - $(\varepsilon,\delta)$-Domain).
Let $0<q,p,\alpha<s<\infty$, let $\Omega\subseteq[0,1]^{d}$ be an $(\epsilon,\delta)$-domain. For each $f\in B^{s}_{p,q}([0,1]^{d})$, every “simultaneous approximation error” $\varepsilon>0$, there exists a Res-KAN $\hat{f}_{\varepsilon}:\mathbb{R}^{d}\to\mathbb{R}$ satisfying
$\big{\|}f-\hat{f}_{\varepsilon}\big{\|}_{B^{\alpha}_{p,q}(\Omega)}<\varepsilon.$
Moreover, $\hat{f}$ has width at-most $(2d+1)\big{(}2^{K+1}-2\big{)}$, depth at-most $d+1$, and no more than $\left(\frac{5d^{2}+23d}{2}+1\right)(2^{K+1}-1)$ non-zero parameters; where $K\leq\big{\lceil}\log_{2}\varepsilon^{1/((\alpha-s)q)}-1\big{\rceil}$.
Proof.
As remarked at the start of [38, Chapter 2] any bounded extension operator on special Lipschitz domains implies a bounded extension operator on general $(\varepsilon,\delta)$-domains (by gluing using a partition of unity); thus, [41, Theorem 1.5] implies that there exists a bounded linear operator $\mathcal{E}:B_{p,q}^{\alpha}(\Omega)\to\mathcal{E}:B_{p,q}^{\alpha}(\mathbb{R}^{d})$. Since the restriction operators between $(\epsilon,\delta)$-domains are bounded then we have the following estimate: for any $f\in B_{p,q}^{\alpha}(\Omega)$ and any Res-KAN $\hat{f}$
$\|f-\hat{f}|_{\Omega}\|_{B_{p,q}^{\alpha}(\Omega)}$ $\lesssim\|\mathcal{E}(f)-\mathcal{E}(\hat{f})|_{\Omega}\|_{B_{p,q}^{\alpha}(\mathbb{R}^{d})}$
$\lesssim\|\mathcal{E}(f)|_{[0,1]^{d}}-\mathcal{E}(\hat{f})|_{\Omega}|_{[0,1]^{d}}\|_{B_{p,q}^{\alpha}([0,1]^{d})}$
$=\|\mathcal{E}(f)|_{[0,1]^{d}}-\mathcal{E}(\hat{f})|_{[0,1]^{d}}\|_{B_{p,q}^{\alpha}([0,1]^{d})}$
$=\|\mathcal{E}(f)|_{[0,1]^{d}}-\hat{f}|_{[0,1]^{d}}\|_{B_{p,q}^{\alpha}([0,1]^{d})}$ (27)
where (27) followed by the definition of the restriction extension operators. Since $\mathcal{E}(f)|_{[0,1]^{d}}\in B_{p,q}^{\alpha}([0,1]^{d})$ then we may retroactively pick our Res-KAN $\hat{f}$ as in Proposition (3) to bound the right-hand side above by $\varepsilon$; i.e.
$\|f-\hat{f}|_{\Omega}\|_{B_{p,q}^{\alpha}(\Omega)}\lesssim\|\mathcal{E}(f)|_{[0,1]^{d}}-\hat{f}|_{[0,1]^{d}}\|_{B_{p,q}^{\alpha}([0,1]^{d})}<\varepsilon.$
This completes our proof. ∎
A.1.2 Fractal Domains
Lemma 5 (KAN-Approximation of Besov Functions on Fractal Domains).
Let $0<\alpha<s<\infty$ and $1\leq p,q<\infty$, $d-1<n<d$ and $\mathcal{X}\subseteq[0,1]^{d}$ be an Ahlfors $n$-regular. Then, for every $f\in B^{s}_{p,q}(\mathcal{X})$ and every $\varepsilon>0$ there is a Res-KAN $\hat{f}$ such that: $\hat{f}|_{\mathcal{X}}$ is indeed a well-defined element of $B_{p,q}^{\alpha-(n-d)/p}(\mathcal{X})$ and satisfies
$\|f-\hat{f}|_{\mathcal{X}}\|_{B^{\alpha-(n-d)/p}_{p,q}(\mathcal{X})}<\varepsilon.$
Moreover, $\hat{f}$ has width at-most $(2d+1)\big{(}2^{K+1}-2\big{)}$, depth at-most $d+1$, and no more than $\left(\frac{5d^{2}+23d}{2}+1\right)(2^{K+1}-1)$ non-zero parameters; where $K\leq\big{\lceil}\log_{2}\varepsilon^{1/((\alpha-(n-d)/p-s)q)}-1\big{\rceil}$.
Proof of Lemma 5.
By the Whitney-type extension result in [21, Theorem 6.1], there exists an $\mathcal{E}:B^{s-(n-d)/p}_{p,q}(\mathcal{X})\to B^{s}_{p,q}(\mathbb{R}^{d})$ and a constant $c_{n,d,p,\mathcal{X}}>0$ such that
$\|\mathcal{E}(f)\|_{B^{s}_{p,q}(\mathcal{X})}\leq c_{n,d,p,\mathcal{X}}\,\|f\|_{B^{s-(n-d)/p}(\mathbb{R}^{d})}$ (28)
for each $f\in B^{s-(n-d)/p}_{p,q}(\mathcal{X})$. Moreover, $\mathcal{E}$ is a left-inverse of the restriction map. Therefore, for each $f\in B^{s}_{p,q}(\mathcal{X})$ and every $\varepsilon>0$, Lemma 3 guarantees that there is a Res-KAN $\hat{f}$ satisfying
$\|\mathcal{E}(f)|_{[0,1]^{d}}-\hat{f}|_{[0,1]^{d}}\|_{B^{\alpha}_{p,q}([0,1]^{d})}<\varepsilon.$ (29)
Since the restriction operators from Besov spaces on $\mathbb{R}^{d}$ to $(\varepsilon,\delta)$-domains are bounded linear operators; then there is a constant $C_{p,q,\alpha,d}>0$ such that: for every $g\in B^{\alpha}_{p,q}([0,1]^{d})$, in particular for $g=\mathcal{E}(f)-\hat{f}$, we have
$\|g\|_{B^{\alpha}_{p,q}(\mathbb{R}^{d})}\leq C_{p,q,\alpha,d}\|g|_{[0,1]^{d}}\|_{B^{\alpha}_{p,q}([0,1]^{d})}.$ (30)
By the restriction theorem in [21, Theorem 4.1.], since $\hat{f}\in B_{p,q}^{\alpha}(\mathbb{R}^{d})$ then its restriction $\hat{f}|_{\mathcal{X}}$ is a well-defined element of $B^{\alpha-(n-d)/p}(\mathbb{R}^{d})$. We may therefore, comfortably combine (28) and (30) with (29) to obtain
$\|f-\hat{f}|_{\mathcal{X}}\|_{B^{\alpha-(n-d)/p}(\mathbb{R}^{d})}$ $\leq c_{n,d,p,\mathcal{X}}\,\|\mathcal{E}(f)-\hat{f}\|_{B^{\alpha-(n-d)/p}(\mathbb{R}^{d})}$
$\leq c_{n,d,p,\mathcal{X}}C_{p,q,\alpha,d}\,\|\mathcal{E}(f)|_{[0,1]^{d}}-\hat{f}|_{[0,1]^{d}}\|_{B^{\alpha}_{p,q}([0,1]^{d})}$
$\leq c_{n,d,p,\mathcal{X}}C_{p,q,\alpha,d}\,\varepsilon.$
Relabeling the constant $\varepsilon>0$ as $\varepsilon/(c_{n,d,p,\mathcal{X}}C_{p,q,\alpha,d})$ yields the conclusion. ∎
We now obtain our main approximation theorem.
Proof of Theorem 1.
Together Lemmata 4 and 5 now imply Theorem 1. ∎
Appendix B Proof of Theorem 2
We now prove our main statistical guarantee.
B.1 Definitions Required for the Proof Theorem 2
Our analysis in this section relies on the following dimensions from classical learning theory.
Definition 3 (Growth function, VC-dimension, Shattering).
Let $\mathcal{H}$ denote a class of functions from $\mathcal{X}$ to $\{0,1\}$ (the hypotheses, or the classification rules). For any non-negative integer $m$, we define the growth function of $\mathcal{H}$ as
$\Pi_{\mathcal{H}}(m)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\max_{x_{1},\ldots,x_{m}\in\mathcal{X}}\left|\{(h(x_{1}),\ldots,h(x_{m})):h\in\mathcal{H}\}\right|.$
If |$\left|\{(h(x_{1}),\ldots,h(x_{m})):h\in\mathcal{H}\}\right|=2^{m}$, we say $\mathcal{H}$ shatters the set $\{x_{1},\ldots,x_{m}\}$. The Vapnik-Chervonenkis dimension of $\mathcal{H}$, denoted $\operatorname{VCdim}(\mathcal{H})$, is the size of the largest shattered set, i.e., the largest $m$ such that $\Pi_{\mathcal{H}}(m)=2^{m}$. If there is no largest $m$, we define $\operatorname{VCdim}(\mathcal{H})=\infty$.
Definition 4 (Pseudodimension).
Let $\mathcal{H}$ be a class of functions from $\mathcal{X}$ to $\mathbb{R}$. The pseudodimension of $\mathcal{H}$, written $\operatorname{Pdim}(\mathcal{H})$, is the largest integer $m$ for which there is an $(x_{i})_{i=1}^{m}\oplus(y_{i})_{i=1}^{m}\in\mathcal{X}^{m}\times\mathbb{R}^{m}$ satisfying: for any $(b_{1},\ldots,b_{m})\in\{0,1\}^{m}$ there exists $h\in\mathcal{H}$ such that
$\forall i:h(x_{i})>y_{i}\iff b_{i}=1$
The pseudodimension is not scale sensitive. This is not the case for the $\gamma$-fat shattering dimension defined as follows.
Definition 5 ($\gamma$-Fat Shattering and $\gamma$-Fat Shattering Dimension).
Let $\mathcal{H}\subseteq[0,1]^{\mathcal{X}}$. We say that $\mathcal{H}$ $P_{\gamma}$-shatters a set $X=\{x_{1},\ldots,x_{n}\}\subseteq\mathcal{X}$ if there exist $s_{1},\ldots,s_{n}$ such that, for all $E\subseteq X$, there is an $h\in\mathcal{H}$ satisfying
$\forall x_{i}\in E,\quad h(x_{i})\geq s_{i}+\gamma$
$\forall x_{i}\in X-E,\quad h(x_{i})\leq s_{i}-\gamma$
The fat shattering dimension of $\mathcal{H}$ at scale $\gamma$, denoted by $\operatorname{Pdim}_{\gamma}(\mathcal{H})$, is the size of a largest $P_{\gamma}$-shattered set.
B.2 The Proof
We now prove our main Learning guarantee.
Lemma 6 (Fat-Shattering Dimension Bound for KANs).
Let $\alpha>0$, $d,L,W,I\in\mathbb{N}_{+}$, with $\lceil\alpha\rceil\leq I$. Then, for every $\gamma>0$
$\operatorname{Pdim}_{\gamma}\big{(}\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})\big{)}\in\mathcal{O}\left(L^{2}\,W^{2}\,(I+1-\lceil\alpha\rceil)\big{(}\log\big{(}L\,W^{2}\,(I+1-\lceil\alpha\rceil)\big{)}+L\big{)}\right).$
To prove Lemma 6, we must first recall the following result relating the pseudo-dimension of a set of binary classifiers implemented by the neural network to the VC-dimension of a modification of that class with two extra computational units. We emphasize that the following result holds for general feedforward neural networks (i.e. neural networks given by a connected directed acyclic graph on which a computation is executed on every node, which is neither initial (input node) nor terminal (output node)); see e.g. [25] for a clean formulation.
Proof.
First, the smoothness condition in (6) implies that each KAN-neuron in (2) can be represented as a feedforward network with computational graph
$\hat{G}=(\{z_{0},z_{2}\}\cup\{z_{1:i}\}_{i=0}^{I+1-\lceil\alpha\rceil},\{\{z_{0},z_{1:i}\}_{i=0}^{I+1-\lceil\alpha\rceil}\cup\{z_{1:i},z_{2}\}_{i=0}^{I+1-\lceil\alpha\rceil}\}$
with input node $z_{0}$, output node $z_{2}$, and for each computational node $z_{1:0},\dots,z_{1:I+1-\lceil\alpha\rceil}$ we have
$z_{1:i}=\beta_{i}p_{i}(x)$
where $p_{i}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathcal{N}_{i}$ (viewed as a piecewise polynomial of degree $I+1$) with, of course, no more than $I+1$ breakpoints.
Each fully-connected Res-KAN layer, as defined in (4), can be represented as a feedforward neural network with $2$ layers (including the input and output layers), and at most $2d_{out}(d_{in}+1)(2(I+1-\lceil\alpha\rceil)$ non-zero parameters, and $(2(I+1-\lceil\alpha\rceil)d_{out}$ computational units. Consequently, every $\hat{f}\in\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})$ can be represented as a feedforward neural network at-most $W^{\prime}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}L2d_{out}(d_{in}+1)(2(I+1-\lceil\alpha\rceil)$ non-zero parameters and at-most $L(2(I+1-\lceil\alpha\rceil)d_{out}$ computational units, arranged into at-most $L^{\prime}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}2L$ layers.
Following [3, Theorem 14.1], for every $\hat{f}\in\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})$ let $\tilde{f}:\mathbb{R}^{d}\to\{0,1\}$ defined by modifying the computational graph of $\hat{f}$ as follows: We added one extra input unit and one extra computation unit. This additional computation unit is a $I_{[0,\infty)}$ (heavyside activation function) unit receiving input only from the output unit of $\hat{f}$ and from the new input unit, and it is the output unit of $\tilde{f}$. Let $\mathcal{F}$ consist of all functions constructed in this manner by modifying some $\hat{f}\in\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})$ in this way.
Thus, [5, Theorem 2.1] implies that the VC-dimension of $\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})$ is at-most
$\operatorname{VCdim}(\mathcal{F})\in\mathcal{O}(W^{\prime}L^{\prime}\log W^{\prime}+W^{\prime}(L^{\prime})^{2}).$ (31)
Now [3, Theorem 14.1] implies that the pseudo-dimension $\operatorname{Pdim}(\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R}))$ satisfies
$\operatorname{Pdim}\big{(}\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})\big{)}\lesssim\operatorname{VCdim}(\mathcal{F}).$ (32)
Combining (31) with (32) yields
$\operatorname{Pdim}\big{(}\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})\big{)}$ $\in\mathcal{O}(W^{\prime}L^{\prime}\log W^{\prime}+W^{\prime}(L^{\prime})^{2})$ (33)
$\in\mathcal{O}\left(L^{2}\,W^{2}\,(I+1-\lceil\alpha\rceil)\big{(}\log\big{(}L\,W^{2}\,(I+1-\lceil\alpha\rceil)\big{)}+L\big{)}\right).$
Now, by [3, Theorem 11.13 (i)], for every $\gamma>0$ we have
$\operatorname{Pdim}_{\gamma}\big{(}\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})\big{)}\leq\operatorname{Pdim}\big{(}\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})\big{)}.$ (34)
We obtain our conclusion upon combining (33) with (34). ∎
Consider the hypothesis class $\mathcal{H}_{W,L}^{\alpha,I,p,q}(\mathcal{X})$ consisting of all maps $h_{\hat{f},f}:\mathcal{X}\to\mathbb{R}$ for which there exist some $\hat{f}\in\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})$ and some $f\in B_{p,q}^{\alpha}(\mathcal{X})$ with $\|f\|_{B_{p,q}^{\alpha}(\mathcal{X})}\leq 1$ such that: for every $x\in\mathcal{X}$ we have
$h_{\hat{f},f}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{|}\hat{f}(x)-f(x)\big{|}.$ (35)
We henceforth denote the unit ball in $B_{p,q}^{\alpha}(\mathcal{X})$ by $B_{p,q}^{\alpha}(\mathcal{X})_{1}$. Our next result bounds the fat-shattering dimension of this “regression error” hypothesis class $\mathcal{H}_{W,L}^{\alpha,I,p,q}(\mathcal{X})$ in terms of the number of the: degree, regularity parameters $I$ and $\alpha$, as well as the depth and width $L$ and $W$ of our hypothesis class, together with some added constraints on the regularity of the target function in the Besov space which we are learning.
Lemma 7 (Fat-Shattering Dimension of the “Regression-Error” Class $\mathcal{H}_{W,L}^{\alpha,I,p,q}(\mathcal{X})$).
Suppose that $\mathcal{X}$ is a Lipschitz domain, let $1\leq\tau\leq\infty$, $1\leq p,q\leq\infty$ and $\alpha>\left(d\left(1/p-1/\tau\right)\right)_{+}$ and let $L,I,W\in\mathbb{N}_{+}$. For every probability measure $\mathbb{P}\in\mathcal{P}(\mathcal{X})$ and every $\gamma>0$ we have
$\operatorname{Pdim}_{\gamma}(\mathcal{H}_{W,L}^{\alpha,I,p,q}(\mathcal{X}))\lesssim\log_{2}(1/(8\gamma))^{2}r\,L^{2}\,W^{2}\big{(}\log\big{(}rL\,W^{2}\big{)}+L\big{)}\Big{)}+(8\gamma)^{-d/\alpha}$
where $r\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}I+1-\lceil\alpha\rceil$.
Proof of Lemma 7.
Abbreviate =$\mathcal{H}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathcal{H}_{W,L}^{\alpha,I,p,q}(\mathcal{X})$. For any $A\subseteq C(\mathcal{X})$ and each $\varepsilon>0$ let $\mathcal{N}(\epsilon,A)$ denote the $\epsilon$-covering number of $A$ in $C(\mathcal{X})$ (with the uniform norm). Then, for every $\varepsilon>0$, the definition of $\mathcal{H}$ and the triangle inequality implies that
$\mathcal{N}(\epsilon,\mathcal{H})\leq\mathcal{N}\big{(}\epsilon/2,\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})\big{)}\mathcal{N}\big{(}\epsilon/2,B_{p,q}^{\alpha}(\mathcal{X})_{1}\big{)}.$ (36)
Taking logarithmic across (36) we find that
$\log_{2}(\mathcal{N}(\epsilon,\mathcal{H}))\leq\log_{2}\Big{(}\mathcal{N}\big{(}\epsilon/2,\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})\big{)}\Big{)}+\log_{2}\Big{(}\mathcal{N}\big{(}\epsilon/2,B_{p,q}^{\alpha}(\mathcal{X})_{1}\big{)}\Big{)}.$ (37)
By [43, Theorem 2.7.4], since $1\leq\tau\leq\infty$, $1\leq p,q\leq\infty$ and $\alpha>\left(d\left(1/p-1/\tau\right)\right)_{+}$ then, we may bound $\log_{2}\Big{(}\mathcal{N}\big{(}\epsilon/2,B_{p,q}^{\alpha}(\mathcal{X})\big{)}\Big{)}$ above by $\mathcal{O}\big{(}(\varepsilon/2)^{-d/\alpha}\big{)}$. Consequently, (37) can be further bounded-above by
$\log_{2}(\mathcal{N}(\epsilon,\mathcal{H}))\lesssim\log_{2}\Big{(}\mathcal{N}\big{(}\epsilon/2,\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})\big{)}\Big{)}+\varepsilon^{-d/\alpha}.$ (38)
Now, applying [6, Theorem 2] we may bound the $\log_{2}$-covering number of
$\log_{2}\Big{(}\mathcal{N}\big{(}\epsilon/2,\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})\big{)}\Big{)}$
above by its $c_{2}\varepsilon$-fat shattering dimension; for some absolute constant $c_{2}>0$ and an additional multiplicative factor of $\log_{2}(1/\varepsilon)^{2}$; that is
$\log_{2}\Big{(}\mathcal{N}\big{(}\epsilon/2,\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})\big{)}\Big{)}$ $\lesssim\operatorname{Pdim}_{c_{2}\,\varepsilon/2}\big{(}\operatorname{Res-KAN}_{L,W}^{I,\alpha}(\mathbb{R}^{d},\mathbb{R})\big{)}\log(2/(c_{2}\varepsilon))^{2}$ (39)
$\in\mathcal{O}\Big{(}\log_{2}(1/\varepsilon)^{2}L^{2}\,W^{2}\,(I+1-\lceil\alpha\rceil)$
×$\times\big{(}\log\big{(}L\,W^{2}\,(I+1-\lceil\alpha\rceil)\big{)}+L\big{)}\Big{)}$
where the order-estimate on the right-hand side of (39) follows from Lemma 6. Incorporating (39) into the right-hand side of (38) yields
$\log_{2}(\mathcal{N}(\epsilon,\mathcal{H}))\lesssim\log_{2}(1/\varepsilon)^{2}L^{2}\,W^{2}\,(I+1-\lceil\alpha\rceil)\big{(}\log\big{(}L\,W^{2}\,(I+1-\lceil\alpha\rceil)\big{)}+L\big{)}\Big{)}+\varepsilon^{-d/\alpha}.$ (40)
Now, applying [6, Theorem 2] again, we have that
$\operatorname{Pdim}_{\epsilon/8}(\mathcal{H})\leq\max_{P}\log_{2}(\mathcal{N}(\epsilon,\mathcal{H},\mathcal{L}_{1}(dP)))\lesssim\log_{2}(\mathcal{N}(\epsilon,\mathcal{H}))$ (41)
where $N(\epsilon,\mathcal{H},\mathcal{L}_{1}(dP))$ is the $\epsilon$-covering number of $\mathcal{H}$ with respect to the norm on $\mathbb{E}_{X\sim\mathbb{P}}[\|X\|]$. Upon using (40) to bound the right-hand side of (41) and then relabelling $\gamma=\varepsilon/8$, we deduce our conclusion. ∎
Having bounded the fat-shattering dimension of our “regression error” hypothesis class $\mathcal{H}_{W,L}^{\alpha,I,p,q}(\mathcal{X})$, we may obtain the conclusion of Theorem 2 by appealing to one of the main results of [2].
Proof of Theorem 2.
We again abbreviate =$\mathcal{H}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathcal{H}_{W,L}^{\alpha,I,p,q}(\mathcal{X})$. By [2, Theorem 3.6], we have that: for every error size $\varepsilon>0$ and each failure probability $0<\delta\leq 1$ the following holds
$\mathbb{P}\Biggl{(}\sup_{h_{f,\hat{f}}\in\mathcal{H}}\,\Big{|}\mathbb{E}_{X\sim\mathbb{P}_{X}}(h_{f,\hat{f}}(X))-\frac{1}{N}\sum_{n=1}^{N}\,h_{f,\hat{f}}(X_{n})\Big{|}\leq\varepsilon\Biggr{)}\geq 1-\delta$ (42)
provided that
$N\leq c_{1}\left(\frac{1}{\epsilon^{2}}\left(\operatorname{Pdim}_{\epsilon/32}(\mathcal{H})\ln^{2}\left(\frac{\operatorname{Pdim}_{\epsilon/32}(\mathcal{H})}{\epsilon}\right)+\ln\left(\frac{1}{\delta}\right)\right)\right)$ (43)
for some absolute constant $c_{1}>1$. Consequently, if we set the failure probability, $\delta$, to be
$\delta\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\exp\Bigl{(}-\tfrac{N\,\epsilon^{2}}{c_{1}}+\operatorname{Pdim}_{\epsilon/32}(\mathcal{H})\,\ln^{2}\Big{(}\tfrac{\operatorname{Pdim}_{\epsilon/32}(\mathcal{H})}{\epsilon}\Big{)}\Bigr{)}.$ (44)
Note that, we may use the upper-bound for the $\varepsilon/32$-fat shattering dimension computed in Lemma 7; namely,
$\operatorname{Pdim}_{\epsilon/32}(\mathcal{H})\lesssim\log_{2}(4/\epsilon)^{2}r\,L^{2}\,W^{2}\big{(}\log\big{(}rL\,W^{2}\big{)}+L\big{)}\Big{)}+(4/\epsilon)^{d/\alpha}$ (45)
where, as before, $r\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(I+1-\lceil\alpha\rceil)$. Therefore such that (42) holds if $N\leq N_{\epsilon,\delta}^{\star}$; where
$N_{\epsilon,\delta}^{\star}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\frac{c_{1}}{\epsilon^{2}}\Bigl{[}\,(A+B)\,\ln^{2}\!\bigl{(}\tfrac{A+B}{\epsilon}\bigr{)}\;+\;\ln\!\bigl{(}\tfrac{1}{\delta}\bigr{)}\Bigr{]}$ (46)
$A$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}c\,\log_{2}\Bigl{(}\tfrac{4}{\epsilon}\Bigr{)}^{2}\,r\,L^{2}\,W^{2}\bigl{(}\ln(rLW^{2})+L\bigr{)}\mbox{ and }B\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(4/\epsilon)^{d/\alpha}$ (47)
for some absolute constant $c>0$. Consequently, we have that
$N\in\mathcal{O}\Bigl{(}\epsilon^{-2-d/\alpha}\,(\ln(1/\varepsilon))^{2}+\epsilon^{-2}\ln(1/\delta)\Bigr{)}.$
Upon combining (42) with (44), and observing that
$\mathcal{R}_{\mathbb{P}}(f|\hat{f})=\mathbb{E}_{X\sim\mathbb{P}_{X}}(h_{f,\hat{f}}(X))\mbox{ and }\hat{\mathcal{R}}_{\mathbb{P}}^{N}(f|\hat{f})=\frac{1}{N}\sum_{n=1}^{N}(h_{f,\hat{f}})(X_{n})$
we obtain our conclusion upon relabelling $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}1/c_{1}$. ∎